OpenStudy (anonymous):

Simplify the following expression. 3 × 22 + 8 ÷ 2 - (5 + 2) A. 17 B. 14 C. 10 D. 9 can someone help me please

OpenStudy (anonymous):

3 x 22= 66, 8/2=4 ,5 + 2=7. So its 66+4-7. 66+ 4=70, and 70-7=33

OpenStudy (anonymous):

so da ansa is 33

OpenStudy (logan13):

this is english